# An Investigation of Prompt Variations for Zero-shot LLM-based Rankers

### v1

@@ 4.2 What are the characteristic of the best prompts?
Table 5 in Appendix B reports the optimal prompt components for various ranking methods and datasets. We identify notable patterns as below.
Task Instruction and Output Type: We analyse task instructions and output types together due to their inherent interrelation; the type of output largely depends on the ranking algorithm the task instruction implements. We analyse these separately for each ranking family because instructions and output types vary across ranking families. We do not analyse pairwise here as for this method there are no alternative choices for these wordings.
@@
Tone Words: The inclusion of tone words does not present a clear trend, with the percentage of prompts achieving a higher effectiveness with no tone words or each of the five tone words distributed as follows: 18% (no tone words), 17%, 18%, 20%, 15%, and 12%. This indicates a relatively uniform influence of tone words choice on prompt effectiveness. Furthermore, no consistent patterns were observed regarding the influence of LLM backbones or datasets on the effectiveness related to tone words. Notably, including a tone word in the prompt led to increased effectiveness in 82% of the cases, underscoring the potential benefit of tone words in enhancing prompt performance.
Role Playing: Role playing often leads to the best effectiveness for pointwise and pairwise prompts (80% and 66% of the cases, respectively), it has mixed effects for setwise (ipresent in the best prompt about half of the times), while it is not associated with the best effectiveness for listwise (13%). Overall, 55% of the prompts with highest effectiveness include role playing wording. Role playing was originally only used by one ranking method, RankGPT Sun et al. (2023), highlighting the difficulty in comparing different ranking methods if their prompt variations are not fairly explored.
Evidence Ordering: For pointwise ranking, presenting passage text before the query text is preferred in 86% of top-performing prompts. The preference is less clear in other ranking types: 40% for pairwise; 40% for listwise; 46% for setwise where the best prompts have the passage text before the query text. Considering model backbones, Flan-T5 tends to perform best when presented with passage text before query text (66% of cases); while results are mixed for the other backbone models.
@@ 4.3 Which ranking method is most effective?
Before our analysis of prompt variations, the only comparison of ranking methods across the four families within a consistent setting of datasets and backbone was provided by Zhuang et al. (2023b). According to their results, the best performing rankers were setwise and pairwise (depending on dataset and backbone), followed by listwise and then pointwise, which were distinguishably worse.
@@
Our analysis of prompt variations reveals a somewhat different picture. Consider the top results for each ranking family across datasets and LLM backbone reported in Figure 1. Similarly to previous findings, we also observe that overall pairwise and setwise methods deliver the best results. However, we find that pointwise can be as competitive as these previous methods if instructed with specific prompts. This is the case throughout all datasets and LLM backbones, with the exception of Llama 3. For Llama 3 on DL datasets, in fact, we observe that pointwise significantly underperforms other methods. At the same time, we also observe that there are instances in which pointwise ranking can far exceed other methods: this is the case on DL19 when using the FLanT5-Large and XL backbones (best pointwise nDCG@10 respectively 0.6918 and 0.7010, statistically significantly outperforming the other methods in most cases), and on DL20 when using Mistral-7B (best pointwise 0.6486).
4.4 Are ranking methods stable?
We consider the four ranking families independently, and study the variance of their effectiveness across all prompt variations for that family. Results are displayed in Figure 1. We observe that pointwise methods display the largest variability in effectiveness due to prompt variations, with some prompt variations delivering poor effectiveness. Setwise and pairwise instead do better, displaying lower variability: setwise achieves this across all backbones and datasets investigated, while pairwise displays larger variability in specific conditions, e.g., when using Mistral, and for the COVID dataset when using Llama3.
Two key insights arise: 1) ranking algorithms are susceptible to prompt variations and may exhibit wide-ranging performance variations, and 2) different ranking algorithms have varying sensitivity to prompts, indicating the potential for some to mitigate the impact of these variations. This suggests more comprehensive prompting optimisation, e.g., via self-prompting, may further improve effectiveness across all rankers.
4.5 Does the LLM size matter?
We answer this question by focusing on the FlanT5 models only, which come in three different sizes (first three rows of plots in Figure 1). We observe that in general larger models deliver higher effectiveness and reduced variations across the board. However, the pointwise family represents an exception to this. Improvements are observed when passing from FlanT5-large to FlanT5-XL. However when using FlanT5-XXL we observe both decreased effectiveness and large variance in effectiveness across the prompt variations.
4.6 Does the LLM backbone matter?
We answer this question by comparing three backbones across the results from Figure 1): FlanT5-XXL (11.3B parameters), Mistral-7B and Llama3-8B: the last two are of comparable size, while the first has approximately 40-60% more parameters. We observe that generally Llama3 and FlanT5 outperform Mistral-based rankers. Llama3 and FlanT5 have overall similar effectiveness, though on the COVID dataset Llama3-based rankers consistently outperform those with FlanT5. We also observe that Mistral-based rankers exhibit larger variance in effectiveness due to prompt variations than rankers based on the other backbones.
5 Conclusions
Recent works have shown the promise of using generative LLMs to implement effective zero-shot re-rankers. Although distinctions among various methods primarily arise from the underlying ranking algorithms, further differences emerge based on the characteristics of the prompts employed to implement these algorithms. These characteristics are not associated with the actual ranking algorithm: e.g., they may be wordings related to a role-playing strategy, or tone words. In this paper, we analysed these prompts and mapped prompt wordings into a set of components to form prompt templates that allowed us to better understand the content of these prompts. We then performed a systematic analysis of prompt variations across different types of zero-shot LLM rankers. Our analysis revealed that ranking effectiveness varies considerably across different implementations of prompt components. Optimal prompt wording showed variability depending on the ranking method, dataset, and LLM backbone employed, suggesting that automatic prompt optimization, tailored to specific ranking methods and datasets, may be more effective than manual prompt engineering for optimizing ranking performance.
We make code, runs and analysis available at https://github.com/ielab/zeroshot-rankers-prompt-variations.
6 Limitations
Latency is an important factor affective the deployment of rankers in production. In this paper, we did not consider query latency as one of the dimensions of analysis and comparison because our focus was on varying the prompts to measure ranking effectiveness. Zhuang et al. provided a first insight into comparing latency among the zero-shot LLM rankers we considered. Latency of the method and prompts we considered in our analysis would be similar to that performed there; however we note that some of the prompts we consider are longer than others (e.g. if the role playing component is added to prompts) – longer prompts result in increased query latency.
Overall, our experiments considered 1,248 prompt variations. While this is a large number, many more prompt variations could have been designed and investigated. However, sensibly increasing the number of prompt variations in our experiments would have been infeasible as it would have exceeded our compute budget. The current set of experiments took in excess of 12,400 GPU-hours to execute. In particular, varying prompt for the pairwise ranking methods results in large computational requirements compared to other methods because of the extensive number of pairwise computations (and thus LLM inferences) required by this method to answer a query.
In our analysis we considered instruction-tuned checkpoints from three LLM backbones: FlanT5 (in three different sizes), Mistral and Llama 3. A wider range of LLMs could have been considered, importantly including OpenAI’s GPT models such as GPT3.5 and GPT4. GPT3.5 for example was found more effective than Llama 2 and Vicuna based LLM rankers in a limited set of experiments in previous work Zhuang et al. (2023b). We were however restricted to using non-commercial models because of the high costs involved in performing the large number of experiments we considered with commercial APIs. For example, using the cost estimates from Zhuang et al., executing our experiments for pairwise, listwise and setwise across all considered prompts using GPT3.5 alone would have costed approximately USD \$5,000.
7 Ethical considerations
In our experiments we considered a large number of prompt variations: for pointwise we consider 768 unique prompts; for pairwise 48; for listwise 288; and for setwise 144 – for a total of 1,248 prompts. We ran these prompts across three datasets, and using five different LLMs with up to 11B parameters. The execution of these experiments required a large amount of compute power; we ran experiments across three clusters: one with Nvidia A100 GPUs, another with Nvidia H100 GPUs, and a smaller one with Nvidia A6000 GPUs. Although we only considered zero-shot LLM ranking approaches and thus did not conduct any LLM training, we recognize that our experiments may have still consumed substantial energy, thereby contributing to CO2 emissions Scells et al. (2022) and water consumption Zuccon et al. (2023).
The publicly available LLMs we used may have been trained with content that contains several types of societal biases Gallegos et al. (2024) – and these biases may permeate in the rankings our zero-shot LLM rankers produced. Future research could explore ways to mitigate these biases prompt engineering.
Appendix A Original Prompts of LLM rankers
Table 4: The content of each original prompt is structured with specific placeholders enclosed in braces “{}”. For instance, “{query}” represents the query text.
Rankers Original Prompt 1 Original Prompt 2 Original Prompt 3 Original Prompt 4
Pointwise
Query: {query}
Passage: {text}
Does the passage answer the query?
Answer ’Yes’ or ’No’
Passage: {text}
Query: {query}
Is this passage relevant to the query?
Please answer True/False. Answer:
For the following query and document,
judge whether they are ’Highly Relevant’, ’Somewhat Relevant’, or ’Not Relevant’.
Query: {query}
Document:{text}
Output:
From a scale of 0 to 4,
judge the relevance between the query and the document.
Query: {query}
Document:{text}
Output:
Pairwise
Given a query: {query}, which of the following two passages is more relevant to the query?
Passage A: {doc1}
Passage B: {doc2}
Output Passage A or Passage B:
Listwise
{Several Passages}
Query = {query}
Passages = [Passage 1, Passage2, Passage3, Passage4]
Sort the Passages by their relevance to the Query.
Sorted Passages = [
You are RankGPT, an intelligent assistant that can rank passages based on their relevancy to the query.
I will provide you with {num} passages, each indicated by number identifier [].
Rank the passages based on their relevance to query: {query}
{Several Passages}
Search Query: {query}.
Rank the {num} passages above based on their relevance to the search query.
The passages should be listed in descending order using identifiers.
The most relevant passages should be listed first. The output format should be [] $>$ [], e.g., [1] $>$ [2].
Only response the ranking results, do not say any word or explain.
Setwise
Given a query "{query}", which of the following passages is the most relevant one to the query?
Passages [Passage A, Passage B, Passage C]
Output only the passage label of the most relevant passage:
The original prompts associated to each LLM ranking method are reported in Table 4; these prompts were collected from the works of Zhuang et al. (2023a); Qin et al. (2023); Ma et al. (2023); Sun et al. (2023); Tang et al. (2023); Zhuang et al. (2023b),.
In the original prompts for listwise the number of passages included in a single prompt is a parameter of the method. All passages included are denoted by “{Several Passages}”, and “{num}” indicates the actual number of passages. Furthermore, in the first original listwise prompt, there is an instruction specifying the format as follows: “Passages = [Passage 1, Passage 2, Passage 3, Passage 4]”. It is important to note that these entries are placeholders, not actual passages, and their count is determined by “{num}”.
Appendix B Best Prompt Variations
Table 5 reports the template combinations that lead to the best performing prompts across combinations of ranking method, LLM backbones, and dataset.
Table 5: Prompt templates containing the components that lead to the most effective results across combinations of ranking method, LLM backbones, and dataset.
@@ Appendix B Best Prompt Variations
Appendix C Effectiveness of Original and Best Prompts
Table 6 reports the nDCG@10 obtained by the original prompts for each ranking method and the best prompt variation for each ranking family.
Table 6: Comparison of nDCG@10 across different rankers, LLMs and datasets. Po: pointwise, Pa: pairwise, Li: listwise, Se: setwise. Original: the original (adapted) prompt. Best: the best prompt found in our experiments. For each dataset and backbone, we highlighted in italics the best performing method that used the original prompt, and in bold the best performing method overall.
Mistral-7B Llama 3-8B
DL19 DL20 COVID DL19 DL20 COVID
Po Qin et al. (2023) 0.5851 0.6699 0.5237 0.6486 0.6897 0.7988 0.4176 0.5328 0.3383 0.4950 0.6678 0.7916
Po Ma et al. (2023) 0.6254 0.5948 0.7379 0.4808 0.3794 0.7527
Po Zhuang et al. (2023a).1 0.6414 0.5951 0.7758 0.3764 0.3088 0.7485
Po Zhuang et al. (2023a).2 0.5577 0.4896 0.6967 0.3890 0.3466 0.7242
Pa Qin et al. (2023) 0.6263 0.6450 0.6036 0.6059 0.7724 0.7738 0.6738 0.6738 0.6163 0.6340 0.7966 0.8012
Li Ma et al. (2023) 0.5863 0.6686 0.5492 0.6348 0.7101 0.7758 0.6364 0.7004 0.6088 0.6619 0.7393 0.8123
Li Sun et al. (2023) 0.5402 0.5184 0.6204 0.5926 0.5377 0.6589
Se Zhuang et al. (2023b) 0.6567 0.6811 0.6180 0.6256 0.7846 0.8053 0.6733 0.7000 0.5990 0.6486 0.7897 0.8035
FlanT5-Large FlanT5-XL FlanT5-XXL
DL19 DL20 COVID DL19 DL20 COVID DL19 DL20 COVID
Po Qin et al. (2023) 0.6180 0.6918 0.5984 0.6266 0.6490 0.7570 0.6157 0.7010 0.6439 0.6727 0.6786 0.7742 0.6289 0.6860 0.6514 0.6733 0.6642 0.7784
Po Ma et al. (2023) 0.6483 0.6128 0.6971 0.6690 0.6426 0.7301 0.6563 0.6533 0.7321
Po Zhuang et al. (2023a).1 0.6468 0.5698 0.6672 0.6300 0.6444 0.7023 0.6451 0.6274 0.6983
Po Zhuang et al. (2023a).2 0.5164 0.4387 0.6206 0.6054 0.5971 0.7095 0.4068 0.3301 0.4335
Pa Qin et al. (2023) 0.6677 0.6724 0.6237 0.6382 0.7558 0.7788 0.6845 0.6986 0.6766 0.6823 0.7536 0.7750 0.6915 0.7135 0.6992 0.7126 0.7452 0.7917
Li Ma et al. (2023) 0.5465 0.6443 0.5081 0.6125 0.6067 0.7521 0.5688 0.6441 0.5576 0.6300 0.6312 0.7438 0.5915 0.7011 0.5852 0.6897 0.6955 0.7834
Li Sun et al. (2023) 0.6199 0.5552 0.7445 0.6129 0.5966 0.6938 0.6920 0.6672 0.7736
Se Zhuang et al. (2023b) 0.6503 0.6693 0.5754 0.6525 0.7440 0.7932 0.6812 0.6959 0.6747 0.6855 0.7540 0.7745 0.6925 0.7047 0.6776 0.7036 0.7617 0.7890

### v2

@@ 4.2 What are the characteristic of the best prompts?
Table 5 reports the optimal prompt components for various ranking methods and datasets. We identify notable patterns as below.
Task Instruction and Output Type: We analyse task instructions (TI) and output types (OT) together due to their inherent interrelation; OT largely depends on the ranking algorithm that the TI implements. We analyse these separately for each ranking family because instructions and output types vary across ranking families. We do not analyse pairwise here as for this method there are no alternative choices for these wordings.
@@
Tone Words: The inclusion of tone words does not present a clear trend, with the percentage of prompts achieving a higher effectiveness with no tone words or each of the five tone words distributed as follows: 18% (no tone words), 17%, 18%, 20%, 15%, and 12%. This indicates a relatively uniform influence of tone words choice on prompt effectiveness. Furthermore, no consistent patterns were observed regarding the influence of LLM backbones or datasets on the effectiveness related to tone words. Nevertheless, including a tone word in the prompt led to increased effectiveness in 82% of the cases, underscoring the potential benefit of tone words in enhancing prompt performance.
Role Playing: Role playing often leads to the best effectiveness for pointwise and pairwise prompts (80% and 66% of the cases, respectively), it has mixed effects for setwise (is present in the best prompt about half of the times), while it is not associated with the best effectiveness for listwise (13%). Overall, 55% of the prompts with highest effectiveness include role playing wording. Role playing was originally only used by one ranking method, RankGPT [22], highlighting the difficulty in comparing different ranking methods if their prompt variations are not fairly explored.
Evidence Ordering: For pointwise ranking, presenting passage text before query text is preferred in 86% of top-performing prompts. The preference is less clear in other ranking types: 40% for pairwise; 40% for listwise; 46% for setwise where the best prompts have the passage text before the query text. Considering model backbones, Flan-T5 tends to perform best when presented with passage text before query text (66% of cases); while results are mixed for the other backbone models.
@@ 4.3 Which ranking method is most effective?
Before our analysis of prompt variations, the only comparison of ranking methods across the four families within a consistent setting of datasets and backbone was provided by Zhuang et al. [34]. According to their results, the best performing rankers were setwise and pairwise (depending on dataset and backbone), followed by listwise and then pointwise, which were distinguishably worse.
@@
Our analysis of prompt variations reveals a somewhat different picture. Consider the top results for each ranking family across datasets and LLM backbone (Figure 1). Similarly to previous findings, we also observe that overall pairwise and setwise methods deliver the best results. However, we find that pointwise can be as competitive as these previous methods if instructed with specific prompts. This is the case throughout all datasets and LLM backbones, with the exception of Llama 3. For Llama 3 on DL datasets, we observe that pointwise significantly underperforms other methods. We also observe that there are instances in which pointwise ranking can far exceed others: this is the case on DL19 when using FLanT5-Large and XL backbones (best pointwise nDCG@10 respectively 0.6918 and 0.7010, in most cases statistically significantly outperforming the others), and on DL20 when using Mistral-7B (best pointwise 0.6486).
Table 5: Prompt templates containing the components that lead to the most effective results across combinations of ranking method, LLM backbones, and dataset. The notation used corresponds to component definitions in Section 3.1 and options in Table 2 (e.g., TI_3 for Task Instruction with the Third Option).
@@
4.4 Are ranking methods stable?
We consider the four ranking families independently, and study the variance of their effectiveness across all prompt variations for that family. Results are displayed in Figure 1. We observe that pointwise methods display the largest variability in effectiveness due to prompt variations, with some prompt variations delivering poor effectiveness. Setwise and pairwise instead do better, displaying lower variability: setwise achieves this across all backbones and datasets investigated, while pairwise displays larger variability in specific conditions, e.g., when using Mistral, and for COVID when using Llama3.
Two key insights arise: 1) ranking algorithms are susceptible to prompt variations and may exhibit wide-ranging performance variations, and 2) different ranking algorithms have varying sensitivity to prompts, indicating the potential for some to mitigate the impact of these variations. This suggests that more comprehensive prompting optimisation, e.g., via self-prompting, may further improve effectiveness across all rankers.
4.5 Does the LLM size matter?
We answer this question by focusing on the FlanT5 models only, which come in three different sizes (first three rows of plots in Figure 1). We observe that in general larger models deliver higher effectiveness and reduced variations across the board. However, the pointwise family represents an exception. Improvements are observed when passing from FlanT5-large to FlanT5-XL. However when using FlanT5-XXL we observe both decreased effectiveness and large variance in effectiveness across the prompt variations.
4.6 Does the LLM backbone matter?
We compare three backbones across the results from Figure 1): FlanT5-XXL (11.3B parameters), Mistral-7B and Llama3-8B: the last two are of comparable size, while the first has approximately 40-60% more parameters. We observe that generally Llama3 and FlanT5 outperform Mistral-based rankers. Llama3 and FlanT5 have overall similar effectiveness, though on COVID Llama3-based rankers consistently outperform those with FlanT5. We also observe that Mistral-based rankers exhibit larger variance in effectiveness due to prompt variations than rankers based on the other backbones.
5 Further Discussion and Limitations
Latency is an important factor affecting the deployment of rankers in production. We did not consider query latency as one of the dimensions of analysis and comparison because our focus was on varying the prompts to measure ranking effectiveness. Zhuang et al. [34] provided a first insight into comparing latency among the zero-shot LLM rankers we considered. Latency of the methods and prompts we considered in our analysis would be similar to that performed there; however we note that some of the prompts we consider are longer than others (e.g., if the role playing component is added to prompts) – longer prompts result in increased query latency.
Overall, our experiments considered 1,248 prompt variations. While this is a large number, many more prompt variations could have been designed and investigated (and the space of prompt variations is virtually infinite). However, sensibly increasing the number of prompt variations in our experiments would have been infeasible as it would have exceeded our compute budget. We ran these prompts across three datasets, and using five different LLMs with up to 11B parameters. The execution of these experiments required a large amount of compute power; we ran experiments across three clusters: one with Nvidia A100 GPUs, another with Nvidia H100 GPUs, and a smaller one with Nvidia A6000 GPUs. Overall, the experiments took in excess of 12,400 GPU-hours to execute. Varying prompt for the pairwise ranking results in large compute requirements compared to other methods because of the extensive number of pairwise computations (and thus LLM inferences) required by this method to answer a query. Although we only considered zero-shot LLM ranking approaches and thus did not conduct any LLM training, we recognize that our experiments may have still consumed substantial energy, thereby contributing to CO2 emissions [21] and water consumption [36].
In our analysis we considered instruction-tuned checkpoints from three LLM backbones. A wider range of LLMs could have been considered, importantly including OpenAI’s GPT models such as GPT3.5 and GPT4. GPT3.5 for example was found more effective than Llama 2 and Vicuna based LLM rankers in a limited set of experiments in previous work [34]. We were however restricted to using non-commercial models because of the high costs involved in performing our large number of experiments with commercial APIs. For example, using the cost estimates from [34], executing our experiments for pairwise, listwise and setwise across all considered prompts using GPT3.5 alone would have costed approximately USD \$5,000.
6 Conclusions
Recent works have shown the promise of using generative LLMs to implement effective zero-shot re-rankers. Although distinctions among various methods primarily arise from the underlying ranking algorithms, further differences emerge based on the characteristics of the prompts employed to implement these algorithms. These characteristics are not associated with the actual ranking algorithm: e.g., they may be wordings related to a role-playing strategy, or tone words. In this paper, we analysed these prompts and mapped prompt wordings into a set of components to form prompt templates that allowed us to better understand the content of these prompts. We then performed a systematic analysis of prompt variations across different types of zero-shot LLM rankers. Our analysis revealed that ranking effectiveness varies considerably across different implementations of prompt components. Optimal prompt wording showed variability depending on the ranking method, dataset, and LLM backbone employed, suggesting that automatic prompt optimization, tailored to specific ranking methods and datasets, may be more effective than manual prompt engineering for optimizing ranking performance. We make code, runs and analysis available at https://github.com/ielab/zeroshot-rankers-prompt-variations.
6.0.1 Acknowledgment.
Shuoqi Sun is partially supported by the Australian Research Council through the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S, CE200100005).